# Leveraging Soft Prompts for Privacy Attacks in Federated Prompt Tuning

### v1

@@ A.6 Proofs
We aim to bound the False Positive Rate, $FPR=\operatorname{Pr}(b^{\prime}=1\mid b=0)$.
The condition $b=0$ means the client has a batch $\mathcal{B}=\{x_{1},\dots,x_{B}\}$ of non-member samples. The adversary guesses $b^{\prime}=1$ if their single adversarial key $k_{a}$ is selected at least once. Let $E_{j}$ be the event that the $j$-th sample $x_{j}\in\mathcal{B}$ ”flips”, by selecting $k_{a}$ as its top-1 nearest neighbor:
$E_{j}=\left\{\|q(x_{j})-k_{a}\|^{2}<\min_{i}\|q(x_{j})-k_{b,i}\|^{2}\right\}$
The total FPR is the probability that at least one of these events occurs (that any sample in the batch flips): $FPR=\operatorname{Pr}(E_{1}\cup E_{2}\cup\dots\cup E_{B})$.
By Boole’s inequality (the Union Bound), we have:
$FPR=\operatorname{Pr}\left(\bigcup_{j=1}^{B}E_{j}\right)\leq\sum_{j=1}^{B}\operatorname{Pr}(E_{j})$
We decompose this sum by the M1M-1 benign clusters (per Assumption 1) from which the $B$ samples are drawn. We have:
$\mu_{FP}=\sum_{j=1}^{B}\operatorname{Pr}(E_{j})=\sum_{i=1}^{M-1}\sum_{j\text{ s.t. }x_{j}\in\text{cluster }i}\operatorname{Pr}(E_{j})$
Since all points from cluster $i$ share the same flip probability, $p_{i}(k_{a})=\operatorname{Pr}(E_{j}\mid x_{j}\in\text{cluster }i)$, the sum simplifies to:
$\mu_{FP}=\sum_{i=1}^{M-1}b_{i}\cdot p_{i}(k_{a})$
The event $E_{j}$ (for a point $x_{j}$ in cluster $i$) requires $k_{a}$ to be closer than all benign keys, which must include its own centroid $k_{b,i}$. Therefore, the probability $p_{i}(k_{a})$ is bounded by the pairwise probability from Lemma 1:
$p_{i}(k_{a})=\operatorname{Pr}\left(\|q(x_{j})-k_{a}\|^{2}<\min_{i^{\prime}}\|q(x_{j})-k_{b,i^{\prime}}\|^{2}\right)\leq\operatorname{Pr}\left(\|q(x_{j})-k_{a}\|^{2}<\|q(x_{j})-k_{b,i}\|^{2}\right)$
Using Lemma 1, we have:
$p_{i}(k_{a})\leq\Phi\left(-\frac{\|k_{b,i}-k_{a}\|}{2\sigma_{i}}\right)$
Now, we find the worst case (highest probability) for this bound. The function $\Phi(-z)$ is maximized when $z$ is minimized. This occurs when the distance $\|k_{b,i}-k_{a}\|$ is minimized, given $k_{a}\in S$. We defined this as $D_{min}(i)=\min_{c\in S}\|k_{b,i}-c\|$. Thus, for any random $k_{a}\in S$, the FPR is bounded by:
$FPR\leq\mu_{FP}=\sum_{i=1}^{M-1}b_{i}\cdot p_{i}(k_{a})\leq\sum_{i=1}^{M-1}b_{i}\Phi\left(-\frac{D_{min}(i)}{2\sigma_{i}}\right)$
A.7 Advantage and Attack Success Rate of Individual models
Figures 911 present a detailed comparison between our proposed PromptMIA and a naive membership inference baseline across three backbone models (ViT-B/32, ConViT, and DeiT) on all four datasets (CIFAR10, CIFAR100, TinyImageNet, and FourDataset). Each subplot reports both the Advantage and the Attack Success Rate (ASR) as a function of the client batch size. PromptMIA performs worse on FourDataset compared to other dataset. This is also the dataset with the lowest predictive accuracy under no DP (see Table 2).
@@ A.7 Advantage and Attack Success Rate of Individual models
A.8 Detailed results on outlier detection
Table 3 reports the full precision, recall, and F1 scores of classical anomaly detection methods applied to the task of detecting adversarial keys in the global prompt pool across all datasets and backbone models. We observe that IsolationForest achieves the highest recall ($\approx 1.0$) on every setting, meaning it successfully flags almost all injected adversarial keys. However, its precision is low (typically 0.180.180.370.37), indicating that many benign keys are incorrectly labeled as adversarial, leading to high false positives. OneClassSVM shows slightly better precision ($\sim 0.15$$0.30$) but still suffers from moderate recall and poor F1 scores. LocalOutlierFactor and EllipticEnvelope fail almost entirely in this scenario, often yielding zero detection or near-zero scores. Moreover, these methods consistently misclassify benign keys as adversarial even when no adversarial keys are present ( see Fig. 12).
@@ A.9 Hyperparameter Analysis
Global Prompt Pool size $M$: Increasing the pool size strengthens the attack. Larger $M$ (e.g., $M=24$) yields consistently higher attack success rates across all batch sizes, while smaller pools (e.g., $M=12$) slightly weaken the attack as batch size grows. When the pool size decreases, the probability of the adversarial keys being selected when $\mathcal{T}\notin\mathcal{D}$ increases, reducing FPR. Since the server is in control of the training protocol and the global prompt pool, they can choose the value $M$. See Fig. 13(a).
@@
Impact of number of training rounds: PromptMIA much higher ASR against models that have been trained for a few rounds compared to randomly initialized keys. This is reflected in our theorectical findings in Section 3.2.3 that FPR is lower when the benign data forms tight and compact clusters in the query space around the benign keys, which happens naturally during the training process. See Fig. 13(f).
@@
A.10 Training Dynamics of Benign Keys
@@
A.11 Membership Inference under very large batch sizes
To assess whether further increasing the batch size can mitigate membership leakage, we conducted an additional set of experiments using an extreme configuration with batch size set to 1024.. As seen in Fig. 15, the attack success rate remains close to 80% on FourDataset and more than 80% on others. We also note that using such batch size is often not possible in practical scenarios where low-resource edge devices cannot afford high VRAM consumption. Defense using large batch size is therefore ineffective and impractical.
@@ A.11 Membership Inference under very large batch sizes
A.12 Performance of PromptMIA against text and multimodal data.
To demonstrate that PromptMIA extends beyond Vision Transformer, we conduct additional experiments on the UPMC Food-101 dataset Gallo et al. (2020) which is a multimodal image-text benchmark containing image–caption pairs. For these experiments, we use the pretrained Vision-and-Language Transformer (ViLT) with a frozen image encoder and frozen LLM-based text encoder (i.e., BERT). We additionally evaluate a text-only configuration by providing only textual inputs. In both the multimodal and text-only cases, PromptMIA achieves strong attack success rates, confirming that the attack is not restricted to the vision domain (see Fig. 16).
@@
A.13 Performance under non-heterogenous settings
@@
A.14 Broader discussion of defense strategies
In this section, we provide in-depth discussion of other potential defense mechanisms against PromptMIA.
  • Secure Aggregation. Another strategy is to apply secure aggregation Bonawitz et al. (2017) so that individual client updates are hidden from the server. The intuition is that, if the server only observes an aggregated result rather than each client’s prompt updates, it may be unable to determine whether a particular adversarial prompt was selected. However, current secure aggregation protocols are developed for linear aggregation (that is, weighted averaging) of local model parameters, and do not extend to probablistic aggregation methods commonly used in federated prompt tuning. Weighted averaging is not sufficient in non-IID settings where local parameters must first be aligned before aggregation in order to prevent important features from collapsing into less informative representations due to semantic misalignment Weng et al. (2024). To address this, the recent PFPT work Weng et al. (2024) developed an aligned and aggregated mechanism based on probabilistic non-parametric clustering which was shown to achieve substantially better performance than weighted averaging in non-IID settings. However, both the alignment and the aggregation steps in PFPT are inherently non-linear, making it unclear how existing secure aggregation methods could be generalized to support such a mechanism.
  • Prompt Dropout. Since PromptMIA relies on all adversarial keys being selected, introducing randomness into the prompt selection (e.g via prompt dropout) can reduce the attack success rate but it remains substantial (larger than 75%) as shown in our experiment with prompt dropout below.
Figure 20: Impact of Prompt Dropout on attack success rate.
A.15 The Use of Large Language Models (LLMs)
Parts of this manuscript were edited with the assistance of a large language model (LLM) to enhance clarity, grammar, and overall readability. The LLM was used solely for language refinement; all ideas, analyses, and results remain entirely the authors’ own.

### v2

@@ A.7 Proofs
With batch size of $1$, the client has a single non-member sample $x$. The FPR is the probability of the event $E_{FP}$ that $q(x)$ selects all $N$ adversarial keys.
$FPR=\Pr(E_{FP})=\Pr\left(\min_{j=1...N}q(x)^{T}k_{a_{j}}>\max_{l=1...M-N}q(x)^{T}k_{b_{l}}\right)$
Let $p_{i}$ be the prior probability that a single non-member sample $x$ belongs to benign cluster $i$ (such that $\sum_{i=1}^{M-N}p_{i}=1$). We find this probability by summing over all MNM-N benign clusters that $q(x)$ could be drawn from:
$FPR=\sum_{i=1}^{M-N}\Pr(E_{FP}\mid q(x)\in\text{cluster }i)\cdot p_{i}$
where $p_{i}=\Pr(q(x)\in\text{cluster }i)$ is the prior probability for cluster $i$.
The term $\Pr(E_{FP}\mid q(x)\in\text{cluster }i)$ is exactly the probability $P(E_{i})$ in Lemma 3.2. We substitute its bound:
$\Pr(E_{FP}\mid i)\leq\min_{\begin{subarray}{c}1\leq j\leq N\\ 1\leq l\leq M-N\end{subarray}}\Phi\left(z_{ijl}(\mathcal{K}_{ADV})\right)$
We place this bound into the above sum:
$FPR\leq\sum_{i=1}^{M-N}p_{i}\cdot\left(\min_{\begin{subarray}{c}1\leq j\leq N\\ 1\leq l\leq M-N\end{subarray}}\Phi\left(z_{ijl}(\mathcal{K}_{ADV})\right)\right)\leq\max_{1\leq i\leq M-N}\left(\min_{\begin{subarray}{c}1\leq j\leq N\\ 1\leq l\leq M-N\end{subarray}}\Phi\left(z_{ijl}(\mathcal{K}_{ADV})\right)\right)$
A.8 Advantage and Attack Success Rate of Individual models
Figures 1012 present a detailed comparison between our proposed PromptMIA and a naive membership inference baseline across three backbone models (ViT-B/32, ConViT, and DeiT) on all four datasets (CIFAR10, CIFAR100, TinyImageNet, and FourDataset). Each subplot reports both the Advantage and the Attack Success Rate (ASR) as a function of the client batch size. PromptMIA performs worse on FourDataset compared to other dataset. This is also the dataset with the lowest predictive accuracy under no DP (see Table 2). Results are averaged across 5 runs.
@@ A.8 Advantage and Attack Success Rate of Individual models
A.9 Detailed results on outlier detection
Table 3 reports the full precision, recall, and F1 scores of classical anomaly detection methods applied to the task of detecting adversarial keys in the global prompt pool across all datasets and backbone models. We observe that IsolationForest achieves the highest recall ($\approx 1.0$) on every setting, meaning it successfully flags almost all injected adversarial keys. However, its precision is low (typically 0.180.180.370.37), indicating that many benign keys are incorrectly labeled as adversarial, leading to high false positives. OneClassSVM shows slightly better precision ($\sim 0.15$$0.30$) but still suffers from moderate recall and poor F1 scores. LocalOutlierFactor and EllipticEnvelope fail almost entirely in this scenario, often yielding zero detection or near-zero scores. Visualization of these outlier detecion methods are given in Fig. 5. Moreover, these methods consistently misclassify benign keys as adversarial even when no adversarial keys are present ( see Fig. 13).
@@ A.10 Hyperparameter Analysis
Global Prompt Pool size $M$: Increasing the pool size strengthens the attack. Larger $M$ (e.g., $M=24$) yields consistently higher attack success rates across all batch sizes, while smaller pools (e.g., $M=12$) slightly weaken the attack as batch size grows. When the pool size increases, the probability of the adversarial keys being selected when $\mathcal{T}\notin\mathcal{D}$ decreases, increasing FPR. Since the server is in control of the training protocol and the global prompt pool, they can choose the value $M$. See Fig. 14(a).
@@
Impact of number of training rounds: PromptMIA much higher ASR against models that have been trained for a few rounds compared to randomly initialized keys. This is reflected in our theorectical findings in Section 3.2.3 that FPR is lower when the benign data forms tight and compact clusters in the query space around the benign keys, which happens naturally during the training process. See Fig. 14(f) and Fig. 15.
@@
A.11 Training Dynamics of Benign Keys
@@
A.12 Membership Inference under very large batch sizes
To assess whether further increasing the batch size can mitigate membership leakage, we conducted an additional set of experiments using an extreme configuration with batch size set to 1024. As seen in Fig. 16, the attack success rate remains close to 80% on FourDataset and more than 80% on others. We also note that using such batch size is often not possible in practical scenarios where low-resource edge devices cannot afford high VRAM consumption. Defense using large batch size is therefore ineffective and impractical.
@@ A.12 Membership Inference under very large batch sizes
A.13 Performance of PromptMIA against text and multimodal data.
To demonstrate that PromptMIA extends beyond Vision Transformer, we conduct additional experiments on the UPMC Food-101 dataset (Gallo et al., 2020) which is a multimodal image-text benchmark containing image–caption pairs. For these experiments, we use the pretrained Vision-and-Language Transformer (ViLT) with a frozen image encoder and frozen LLM-based text encoder (i.e., BERT). We additionally evaluate a text-only configuration by providing only textual inputs. In both the multimodal and text-only cases, PromptMIA achieves strong attack success rates, confirming that the attack is not restricted to the vision domain (see Fig. 17).
@@
A.14 Performance under non-heterogenous settings